Sketch the region enclosed by the given curves and find its area. y = 3 – x2 and y = x2 – 15.

Help me this homework please, pls detail as much as possible
Answer #1

५% - +3 = 2-15 24-15 (13) (-3,0) (3,0) (SON (-) (-3,0) (IST) । त for intersection point -2+3= -७ 262- 18 4- 9 ३६% +3 3 -13 2.

- +18 ॥ 2 - 18 + 54] %3D 2x36 = 72

I have solved the problem if you have any problem with the solution then please tell me on the comment box